\begin{thebibliography}{73}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Abdel-Salam and Rafea(2022)]{abdel2022performance}
Shehab Abdel-Salam and Ahmed Rafea.
\newblock Performance study on extractive text summarization using bert models.
\newblock \emph{Information}, 13\penalty0 (2):\penalty0 67, 2022.

\bibitem[Aghajanyan et~al.(2021)Aghajanyan, Gupta, Shrivastava, Chen, Zettlemoyer, and Gupta]{aghajanyan2021muppet}
Armen Aghajanyan, Anchit Gupta, Akshat Shrivastava, Xilun Chen, Luke Zettlemoyer, and Sonal Gupta.
\newblock Muppet: Massive multi-task representations with pre-finetuning.
\newblock In \emph{Proc. EMNLP}, pages 5799--5811, 2021.
\newblock \doi{10.18653/v1/2021.emnlp-main.468}.

\bibitem[Almazrouei et~al.(2023)Almazrouei, Alobeidli, Alshamsi, Cappelli, Cojocaru, Debbah, Goffinet, Heslow, Launay, Malartic, Noune, Pannier, and Penedo]{falcon}
Ebtesam Almazrouei, Hamza Alobeidli, Abdulaziz Alshamsi, Alessandro Cappelli, Ruxandra Cojocaru, Merouane Debbah, Etienne Goffinet, Daniel Heslow, Julien Launay, Quentin Malartic, Badreddine Noune, Baptiste Pannier, and Guilherme Penedo.
\newblock {Falcon-40B}: an open large language model with state-of-the-art performance, 2023.
\newblock Placeholder, paper not yet published.

\bibitem[An et~al.(2021)An, Zhong, Geng, Yang, and Qiu]{an2021retrievalsum}
Chenxin An, Ming Zhong, Zhichao Geng, Jianqiang Yang, and Xipeng Qiu.
\newblock Retrievalsum: A retrieval enhanced framework for abstractive summarization.
\newblock \emph{arXiv preprint arXiv:2109.07943}, 2021.

\bibitem[Bekoulis et~al.(2018)Bekoulis, Deleu, Demeester, and Develder]{bekoulis2018joint}
Giannis Bekoulis, Johannes Deleu, Thomas Demeester, and Chris Develder.
\newblock Joint entity recognition and relation extraction as a multi-head selection problem.
\newblock \emph{Expert Systems with Applications}, 114:\penalty0 34--45, 2018.

\bibitem[Beltagy et~al.(2020)Beltagy, Peters, and Cohan]{beltagy2020longformer}
Iz~Beltagy, Matthew~E. Peters, and Arman Cohan.
\newblock Longformer: The long-document transformer, 2020.

\bibitem[Bhatia et~al.(2019)Bhatia, Celikkaya, Khalilia, and Senthivel]{bhatia2019comprehend}
Parminder Bhatia, Busra Celikkaya, Mohammed Khalilia, and Selvan Senthivel.
\newblock Comprehend medical: a named entity recognition and relationship extraction web service.
\newblock In \emph{ICMLA}, pages 1844--1851. IEEE, 2019.

\bibitem[Bishop(1995)]{bishop1995training}
Chris~M Bishop.
\newblock Training with noise is equivalent to tikhonov regularization.
\newblock \emph{Neural computation}, 7\penalty0 (1):\penalty0 108--116, 1995.
\newblock \doi{10.1162/neco.1995.7.1.108}.

\bibitem[Breiman(1995)]{breiman1995better}
Leo Breiman.
\newblock Better subset regression using the nonnegative garrote.
\newblock \emph{Technometrics}, 37\penalty0 (4):\penalty0 373--384, 1995.
\newblock \doi{10.1080/00401706.1995.10484371}.

\bibitem[Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, et~al.]{gpt3}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et~al.
\newblock Language models are few-shot learners.
\newblock In \emph{Proc. NeurIPS}, volume~33, pages 1877--1901, 2020.
\newblock URL \url{https://proceedings.neurips.cc/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf}.

\bibitem[Chalkidis et~al.(2019)Chalkidis, Androutsopoulos, and Aletras]{chalkidis2019neural}
Ilias Chalkidis, Ion Androutsopoulos, and Nikolaos Aletras.
\newblock Neural legal judgment prediction in english.
\newblock In \emph{Proc. ACL}, pages 4317--4323, 2019.
\newblock \doi{10.18653/v1/P19-1424}.

\bibitem[Chen et~al.(2020)Chen, Sun, Yang, and Lin]{chen2020joint}
Yanguang Chen, Yuanyuan Sun, Zhihao Yang, and Hongfei Lin.
\newblock Joint entity and relation extraction for legal documents with legal feature enhancement.
\newblock In \emph{Proc. COLING}, pages 1561--1571, 2020.

\bibitem[Cho et~al.(2014)Cho, van Merri{\"e}nboer, Bahdanau, and Bengio]{cho2014properties}
Kyunghyun Cho, Bart van Merri{\"e}nboer, Dzmitry Bahdanau, and Yoshua Bengio.
\newblock On the properties of neural machine translation: Encoder--decoder approaches.
\newblock In \emph{Proc. SSST-8}, pages 103--111, 2014.

\bibitem[Clark et~al.(2020)Clark, Luong, Le, and Manning]{Clark2020ELECTRA}
Kevin Clark, Minh-Thang Luong, Quoc~V. Le, and Christopher~D. Manning.
\newblock {ELECTRA}: Pre-training text encoders as discriminators rather than generators.
\newblock In \emph{Proc. ICLR}, 2020.
\newblock \doi{10.48550/arXiv.2003.10555}.

\bibitem[Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek, Guzm{\'a}n, Grave, Ott, Zettlemoyer, and Stoyanov]{conneau-etal-2020-unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov.
\newblock Unsupervised cross-lingual representation learning at scale.
\newblock In \emph{Proc. ACL}, pages 8440--8451, 2020.
\newblock \doi{10.18653/v1/2020.acl-main.747}.

\bibitem[Deu{\ss}er et~al.(2022)Deu{\ss}er, Ali, Hillebrand, Nurchalifah, Jacob, Bauckhage, and Sifa]{deusser2022kpiedgar}
Tobias Deu{\ss}er, Syed~Musharraf Ali, Lars Hillebrand, Desiana Nurchalifah, Basil Jacob, Christian Bauckhage, and Rafet Sifa.
\newblock {KPI-EDGAR}: A novel dataset and accompanying metric for relation extraction from financial documents.
\newblock In \emph{Proc. ICMLA}, pages 1654--1659, 2022.
\newblock \doi{10.1109/ICMLA55696.2022.00254}.

\bibitem[Deu{\ss}er et~al.(2023)Deu{\ss}er, Pielka, Pucknat, Jacob, Dilmaghani, Nourimand, Kliem, Loitz, Bauckhage, and Sifa]{deusser2023contradiction}
Tobias Deu{\ss}er, Maren Pielka, Lisa Pucknat, Basil Jacob, Tim Dilmaghani, Mahdis Nourimand, Bernd Kliem, R{\"u}diger Loitz, Christian Bauckhage, and Rafet Sifa.
\newblock Contradiction detection in financial reports.
\newblock In \emph{Proc. NLDL}, volume~4, 2023.
\newblock \doi{10.7557/18.6799}.

\bibitem[Deußer et~al.(2023)Deußer, Hillebrand, Bauckhage, and Sifa]{deusser2023informed}
Tobias Deußer, Lars Hillebrand, Christian Bauckhage, and Rafet Sifa.
\newblock Informed named entity recognition decoding for generative language models, 2023.

\bibitem[Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova]{devlin2018bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
\newblock {BERT}: Pre-training of deep bidirectional transformers for language understanding.
\newblock In \emph{Proc. NAACL-HLT}, 2019.
\newblock \doi{10.18653/v1/N19-1423}.

\bibitem[Dodge et~al.(2020)Dodge, Ilharco, Schwartz, Farhadi, Hajishirzi, and Smith]{dodge2020finetuning}
Jesse Dodge, Gabriel Ilharco, Roy Schwartz, Ali Farhadi, Hannaneh Hajishirzi, and Noah Smith.
\newblock Fine-tuning pretrained language models: Weight initializations, data orders, and early stopping, 2020.

\bibitem[Fu et~al.(2019)Fu, Li, and Ma]{fu2019graphrel}
Tsu-Jui Fu, Peng-Hsuan Li, and Wei-Yun Ma.
\newblock Graphrel: Modeling text as relational graphs for joint entity and relation extraction.
\newblock In \emph{Proc. ACL}, pages 1409--1418, 2019.

\bibitem[Geng et~al.(2021)Geng, Zhang, and Han]{geng2021joint}
Zhiqiang Geng, Yanhui Zhang, and Yongming Han.
\newblock Joint entity and relation extraction model based on rich semantics.
\newblock \emph{Neurocomputing}, 429:\penalty0 132--140, 2021.

\bibitem[Giorgi et~al.(2019)Giorgi, Wang, Sahar, Shin, Bader, and Wang]{giorgi2019end}
John Giorgi, Xindi Wang, Nicola Sahar, Won~Young Shin, Gary~D Bader, and Bo~Wang.
\newblock End-to-end named entity recognition and relation extraction using pre-trained language models.
\newblock \emph{arXiv preprint arXiv:1912.13415}, 2019.

\bibitem[Girosi et~al.(1995)Girosi, Jones, and Poggio]{girosi1995regularization}
Federico Girosi, Michael Jones, and Tomaso Poggio.
\newblock Regularization theory and neural networks architectures.
\newblock \emph{Neural computation}, 7\penalty0 (2):\penalty0 219--269, 1995.
\newblock \doi{10.1162/neco.1995.7.2.219}.

\bibitem[Gupta et~al.(2016)Gupta, Sch{\"u}tze, and Andrassy]{gupta2016table}
Pankaj Gupta, Hinrich Sch{\"u}tze, and Bernt Andrassy.
\newblock Table filling multi-task recurrent neural network for joint entity and relation extraction.
\newblock In \emph{Proc, COLING}, pages 2537--2547, 2016.

\bibitem[Hanson and Pratt(1988)]{hanson1988comparing}
Stephen Hanson and Lorien Pratt.
\newblock Comparing biases for minimal network construction with back-propagation.
\newblock In D.~Touretzky, editor, \emph{Proc. NeurIPS}, volume~1, 1988.

\bibitem[Hillebrand et~al.(2022)Hillebrand, Deußer, Dilmaghani, Kliem, Loitz, Bauckhage, and Sifa]{kpibert}
Lars Hillebrand, Tobias Deußer, Tim Dilmaghani, Bernd Kliem, Rüdiger Loitz, Christian Bauckhage, and Rafet Sifa.
\newblock {KPI-BERT}: A joint named entity recognition and relation extraction model for financial reports.
\newblock In \emph{Proc. ICPR}, pages 606--612, 2022.
\newblock \doi{10.1109/ICPR56361.2022.9956191}.

\bibitem[Hinton et~al.(2012)Hinton, Srivastava, Krizhevsky, Sutskever, and Salakhutdinov]{hinton2012improving}
Geoffrey~E. Hinton, Nitish Srivastava, Alex Krizhevsky, Ilya Sutskever, and Ruslan~R. Salakhutdinov.
\newblock Improving neural networks by preventing co-adaptation of feature detectors, 2012.

\bibitem[Hochreiter and Schmidhuber(1997)]{lstm}
Sepp Hochreiter and Jürgen Schmidhuber.
\newblock Long short-term memory.
\newblock \emph{Neural Computation}, 9\penalty0 (8):\penalty0 1735--1780, 1997.
\newblock \doi{10.1162/neco.1997.9.8.1735}.

\bibitem[Hoerl and Kennard(1970)]{hoerl1970ridge}
Arthur~E Hoerl and Robert~W Kennard.
\newblock Ridge regression: Biased estimation for nonorthogonal problems.
\newblock \emph{Technometrics}, 12\penalty0 (1):\penalty0 55--67, 1970.
\newblock \doi{10.1080/00401706.1970.10488634}.

\bibitem[Howard and Ruder(2018)]{howard-ruder-2018-universal}
Jeremy Howard and Sebastian Ruder.
\newblock Universal language model fine-tuning for text classification.
\newblock In \emph{Proc. ACL}, pages 328--339, 2018.
\newblock \doi{10.18653/v1/P18-1031}.

\bibitem[Jabbari et~al.(2020)Jabbari, Sauvage, Zeine, and Chergui]{jabbari2020french}
Ali Jabbari, Olivier Sauvage, Hamada Zeine, and Hamza Chergui.
\newblock A french corpus and annotation schema for named entity recognition and relation extraction of financial news.
\newblock In \emph{Proc. LREC}, pages 2293--2299, 2020.

\bibitem[Jain et~al.(2022)Jain, Borah, and Biswas]{jain2022improving}
Deepali Jain, Malaya~Dutta Borah, and Anupam Biswas.
\newblock Improving kullback-leibler based legal document summarization using enhanced text representation.
\newblock In \emph{2022 IEEE SILCON}, pages 1--5. IEEE, 2022.

\bibitem[Jiang et~al.(2020)Jiang, He, Chen, Liu, Gao, and Zhao]{jiang-etal-2020-smart}
Haoming Jiang, Pengcheng He, Weizhu Chen, Xiaodong Liu, Jianfeng Gao, and Tuo Zhao.
\newblock {SMART}: Robust and efficient fine-tuning for pre-trained natural language models through principled regularized optimization.
\newblock In \emph{Proc. ACL}, pages 2177--2190, 2020.
\newblock \doi{10.18653/v1/2020.acl-main.197}.

\bibitem[Kambar et~al.(2022)Kambar, Esmaeilzadeh, and Heidari]{kambar2022survey}
Mina Esmail Zadeh~Nojoo Kambar, Armin Esmaeilzadeh, and Maryam Heidari.
\newblock A survey on deep learning techniques for joint named entities and relation extraction.
\newblock In \emph{IEEE AIIoT}, pages 218--224. IEEE, 2022.

\bibitem[Koniaris et~al.(2023)Koniaris, Galanis, Giannini, and Tsanakas]{koniaris2023evaluation}
Marios Koniaris, Dimitris Galanis, Eugenia Giannini, and Panayiotis Tsanakas.
\newblock Evaluation of automatic legal text summarization techniques for greek case law.
\newblock \emph{Information}, 14\penalty0 (4):\penalty0 250, 2023.

\bibitem[Kornilova and Eidelman(2019)]{kornilova-eidelman-2019-billsum}
Anastassia Kornilova and Vladimir Eidelman.
\newblock {B}ill{S}um: A corpus for automatic summarization of {US} legislation.
\newblock In \emph{Proc. New Frontiers in Summarization}, pages 48--56, 2019.
\newblock \doi{10.18653/v1/D19-5406}.

\bibitem[Lauren\c{c}on et~al.(2022)Lauren\c{c}on, Saulnier, Wang, Akiki, Villanova~del Moral, Le~Scao, Von~Werra, Mou, Gonz\'{a}lez~Ponferrada, Nguyen, et~al.]{laurenccon2022bigscience}
Hugo Lauren\c{c}on, Lucile Saulnier, Thomas Wang, Christopher Akiki, Albert Villanova~del Moral, Teven Le~Scao, Leandro Von~Werra, Chenghao Mou, Eduardo Gonz\'{a}lez~Ponferrada, Huu Nguyen, et~al.
\newblock The bigscience roots corpus: A 1.6tb composite multilingual dataset.
\newblock In \emph{Proc. NeurIPS}, volume~35, pages 31809--31826, 2022.

\bibitem[Lee et~al.(2020)Lee, Cho, and Kang]{Lee2020Mixout}
Cheolhyoung Lee, Kyunghyun Cho, and Wanmo Kang.
\newblock Mixout: Effective regularization to finetune large-scale pretrained language models.
\newblock In \emph{Proc. ICLR}, 2020.
\newblock URL \url{https://openreview.net/forum?id=HkgaETNtDB}.

\bibitem[Li and Ji(2014)]{li2014incremental}
Qi~Li and Heng Ji.
\newblock Incremental joint extraction of entity mentions and relations.
\newblock In \emph{Proc. ACL}, pages 402--412, 2014.

\bibitem[Liang et~al.(2022)Liang, Li, Wu, Zeng, Jiang, Li, and Li]{liang2022efficient}
Xinnian Liang, Jing Li, Shuangzhi Wu, Jiali Zeng, Yufan Jiang, Mu~Li, and Zhoujun Li.
\newblock An efficient coarse-to-fine facet-aware unsupervised summarization framework based on semantic blocks.
\newblock \emph{arXiv preprint arXiv:2208.08253}, 2022.

\bibitem[Lin(2004)]{lin-2004-rouge}
Chin-Yew Lin.
\newblock {ROUGE}: A package for automatic evaluation of summaries.
\newblock In \emph{Text Summarization Branches Out}, pages 74--81, 2004.
\newblock URL \url{https://aclanthology.org/W04-1013}.

\bibitem[Liu and Lapata(2019)]{liu2019text}
Yang Liu and Mirella Lapata.
\newblock Text summarization with pretrained encoders.
\newblock \emph{arXiv preprint arXiv:1908.08345}, 2019.

\bibitem[Liu et~al.(2020)Liu, Shen, and Lapata]{liu2020noisy}
Yang Liu, Sheng Shen, and Mirella Lapata.
\newblock Noisy self-knowledge distillation for text summarization.
\newblock \emph{arXiv preprint arXiv:2009.07032}, 2020.

\bibitem[Loshchilov and Hutter(2019)]{loshchilov2018decoupled}
Ilya Loshchilov and Frank Hutter.
\newblock Decoupled weight decay regularization.
\newblock In \emph{Proc. ICLR}, 2019.
\newblock URL \url{https://openreview.net/forum?id=Bkg6RiCqY7}.

\bibitem[Mahmoud and Hafez(2022)]{mahmoud2022novel}
Hanan A~Hosni Mahmoud and Alaaeldin~M Hafez.
\newblock A novel optimized language-independent text summarization technique.
\newblock \emph{CMC}, 73\penalty0 (3), 2022.

\bibitem[McCloskey and Cohen(1989)]{mccloskey1989catastrophic}
Michael McCloskey and Neal~J Cohen.
\newblock Catastrophic interference in connectionist networks: The sequential learning problem.
\newblock In \emph{Psychology of learning and motivation}, volume~24, pages 109--165. Elsevier, 1989.
\newblock \doi{10.1016/S0079-7421(08)60536-8}.

\bibitem[Merity et~al.(2018)Merity, Keskar, and Socher]{merity2018regularizing}
Stephen Merity, Nitish~Shirish Keskar, and Richard Socher.
\newblock Regularizing and optimizing {LSTM} language models.
\newblock In \emph{Proc. ICLR}, 2018.
\newblock URL \url{https://openreview.net/forum?id=SyyGPP0TZ}.

\bibitem[Miwa and Sasaki(2014)]{miwa2014modelling}
Makoto Miwa and Yutaka Sasaki.
\newblock Modeling joint entity and relation extraction with table representation.
\newblock In \emph{Proc. EMNLP}, pages 1858--1869, 2014.

\bibitem[OpenAI(2023)]{openai2023gpt4}
OpenAI.
\newblock {GPT-4} technical report, 2023.

\bibitem[Ott et~al.(2018)Ott, Auli, Grangier, and Ranzato]{ott2018analyzing}
Myle Ott, Michael Auli, David Grangier, and Marc’Aurelio Ranzato.
\newblock Analyzing uncertainty in neural machine translation.
\newblock In \emph{Proc. ICML}, volume~35, pages 3956--3965. PMLR, 2018.
\newblock URL \url{https://proceedings.mlr.press/v80/ott18a.html}.

\bibitem[Pan et~al.(2023)Pan, Guo, Deng, Yang, Chen, and Chen]{pan2023improving}
Haolin Pan, Yong Guo, Qinyi Deng, Haomin Yang, Jian Chen, and Yiqun Chen.
\newblock Improving fine-tuning of self-supervised models with contrastive initialization.
\newblock \emph{Neural Networks}, 159:\penalty0 198--207, 2023.
\newblock \doi{10.1016/j.neunet.2022.12.012}.

\bibitem[Prechelt(2002)]{prechelt2002early}
Lutz Prechelt.
\newblock Early stopping-but when?
\newblock In \emph{Neural Networks: Tricks of the trade}, pages 55--69. Springer, 2002.

\bibitem[Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu]{t5}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu.
\newblock Exploring the limits of transfer learning with a unified text-to-text transformer.
\newblock \emph{The Journal of Machine Learning Research}, 21\penalty0 (1), jan 2020.
\newblock ISSN 1532-4435.
\newblock \doi{10.48550/arXiv.1910.10683}.

\bibitem[Ramamurthy et~al.(2023)Ramamurthy, Ammanabrolu, Brantley, Hessel, Sifa, Bauckhage, Hajishirzi, and Choi]{ramamurthy2023is}
Rajkumar Ramamurthy, Prithviraj Ammanabrolu, Kiant{\'e} Brantley, Jack Hessel, Rafet Sifa, Christian Bauckhage, Hannaneh Hajishirzi, and Yejin Choi.
\newblock Is reinforcement learning (not) for natural language processing: Benchmarks, baselines, and building blocks for natural language policy optimization.
\newblock In \emph{Proc. ICLR}, 2023.
\newblock URL \url{https://openreview.net/forum?id=8aHzds2uUyB}.

\bibitem[Rehman et~al.(2022)Rehman, Das, Sanyal, and Chattopadhyay]{rehman2022analysis}
Tohida Rehman, Suchandan Das, Debarshi~Kumar Sanyal, and Samiran Chattopadhyay.
\newblock An analysis of abstractive text summarization using pre-trained models.
\newblock In \emph{Proc. IEM-ICDC}, pages 253--264. Springer, 2022.

\bibitem[Scao et~al.(2023)Scao, Fan, Akiki, Pavlick, Ilić, Hesslow, Castagné, Luccioni, Yvon, Gallé, et~al.]{workshop2023bloom}
Teven~Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra~Sasha Luccioni, François Yvon, Matthias Gallé, et~al.
\newblock {BLOOM}: A 176b-parameter open-access multilingual language model, 2023.

\bibitem[Schuster and Nakajima(2012)]{schuster2012japanese}
Mike Schuster and Kaisuke Nakajima.
\newblock Japanese and korean voice search.
\newblock In \emph{Proc. ICASSP}, 2012.

\bibitem[Shang et~al.(2022)Shang, Huang, and Mao]{shang2022onerel}
Yu-Ming Shang, Heyan Huang, and Xianling Mao.
\newblock Onerel: Joint entity and relation extraction with one module in one step.
\newblock In \emph{Proc. AAAI}, pages 11285--11293, 2022.

\bibitem[Sheik and Nirmala(2021)]{sheik2021deep}
Reshma Sheik and S~Jaya Nirmala.
\newblock Deep learning techniques for legal text summarization.
\newblock In \emph{Proc. UP-CON}, pages 1--5. IEEE, 2021.

\bibitem[Sun et~al.(2019)Sun, Qiu, Xu, and Huang]{sun2019fine}
Chi Sun, Xipeng Qiu, Yige Xu, and Xuanjing Huang.
\newblock How to fine-tune {BERT} for text classification?
\newblock In \emph{Proc. CCL}, pages 194--206. Springer, 2019.

\bibitem[Tenney et~al.(2019)Tenney, Das, and Pavlick]{tenney-etal-2019-bert}
Ian Tenney, Dipanjan Das, and Ellie Pavlick.
\newblock {BERT} rediscovers the classical {NLP} pipeline.
\newblock In \emph{Proc. ACL}, pages 4593--4601, 2019.
\newblock \doi{10.18653/v1/P19-1452}.

\bibitem[Touvron et~al.(2023)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et~al.]{touvron2023llama}
Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et~al.
\newblock {LLaMA}: Open and efficient foundation language models, 2023.

\bibitem[Trosten(2023)]{trosten2023questionable}
Daniel~J Trosten.
\newblock Questionable practices in methodological deep learning research.
\newblock In \emph{Proc. NLDL}, volume~4, 2023.
\newblock \doi{10.7557/18.6804}.

\bibitem[Wang et~al.(2021)Wang, Jiang, Bach, Wang, Huang, Huang, and Tu]{wang2021automated}
Xinyu Wang, Yong Jiang, Nguyen Bach, Tao Wang, Zhongqiang Huang, Fei Huang, and Kewei Tu.
\newblock Automated concatenation of embeddings for structured prediction.
\newblock In \emph{Proc. ACL-IJCNLP}, pages 2643--2660, 2021.
\newblock \doi{10.18653/v1/2021.acl-long.206}.

\bibitem[Wu et~al.(2022)Wu, Wu, Qi, and Huang]{wu2022noisytune}
Chuhan Wu, Fangzhao Wu, Tao Qi, and Yongfeng Huang.
\newblock {NoisyTune}: A little noise can help you finetune pretrained language models better.
\newblock In \emph{Proc. ACL}, pages 680--685, 2022.
\newblock \doi{10.18653/v1/2022.acl-short.76}.

\bibitem[Xue et~al.(2019)Xue, Zhou, Ma, Ruan, Zhang, and He]{xue2019fine}
Kui Xue, Yangming Zhou, Zhiyuan Ma, Tong Ruan, Huanhuan Zhang, and Ping He.
\newblock Fine-tuning bert for joint entity and relation extraction in chinese medical text.
\newblock In \emph{BIBM}, pages 892--897. IEEE, 2019.

\bibitem[Ye et~al.(2022)Ye, Lin, Li, and Sun]{ye2022packed}
Deming Ye, Yankai Lin, Peng Li, and Maosong Sun.
\newblock Packed levitated marker for entity and relation extraction.
\newblock In \emph{Proc. ACL}, pages 4904--4917, 2022.
\newblock \doi{10.18653/v1/2022.acl-long.337}.

\bibitem[Yousefi-Azar and Hamey(2017)]{yousefi2017text}
Mahmood Yousefi-Azar and Len Hamey.
\newblock Text summarization using unsupervised deep learning.
\newblock \emph{Expert Systems with Applications}, 68:\penalty0 93--105, 2017.

\bibitem[Yu et~al.(2019)Yu, Zhang, Shu, Wang, Liu, Wang, and Li]{yu2019joint}
Bowen Yu, Zhenyu Zhang, Xiaobo Shu, Yubin Wang, Tingwen Liu, Bin Wang, and Sujian Li.
\newblock Joint extraction of entities and relations based on a novel decomposition strategy.
\newblock \emph{arXiv preprint arXiv:1909.04273}, 2019.

\bibitem[Zhang et~al.(2019)Zhang, Xu, and Wang]{zhang2019pretraining}
Haoyu Zhang, Jianjun Xu, and Ji~Wang.
\newblock Pretraining-based natural language generation for text summarization.
\newblock \emph{arXiv preprint arXiv:1902.09243}, 2019.

\bibitem[Zhang et~al.(2022)Zhang, Roller, Goyal, Artetxe, Chen, Chen, Dewan, Diab, Li, Lin, et~al.]{zhang2022opt}
Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui Chen, Christopher Dewan, Mona Diab, Xian Li, Xi~Victoria Lin, et~al.
\newblock {OPT}: Open pre-trained transformer language models, 2022.

\bibitem[Zheng et~al.(2017)Zheng, Wang, Bao, Hao, Zhou, and Xu]{zheng2017joint}
Suncong Zheng, Feng Wang, Hongyun Bao, Yuexing Hao, Peng Zhou, and Bo~Xu.
\newblock Joint extraction of entities and relations based on a novel tagging scheme.
\newblock \emph{arXiv preprint arXiv:1706.05075}, 2017.

\end{thebibliography}

\documentclass[table]{article}

\usepackage{arxiv}

\usepackage{amsmath}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{cleveref}       
\usepackage{graphicx}
\usepackage[round]{natbib}
\usepackage{doi}

\usepackage{tikz}
\usepackage{tcolorbox}
\usepackage{mathtools}
\usepackage{amsthm}
\usepackage{bbm}
\usepackage{bm}
\usepackage{tabularx}
\usepackage{threeparttable}
\usepackage{makecell}
\usepackage{multirow}
\usepackage{booktabs}
\usepackage{subcaption}
\usepackage{flushend}
\usepackage{dirtytalk}
\usepackage{xcolor}

\definecolor{fhgreen}{RGB}{0,147,116}
\definecolor{orange}{RGB}{	255, 127, 0}
\definecolor{unibonnblue}{HTML}{0a55a1}

\title{Controlled Randomness Improves the Performance of Transformer Models}

\date{}

\usepackage{authblk}

\author[1,2]{%
	Tobias Deu{\ss}er\thanks{\texttt{tdeusser@uni-bonn.de}, ORCID-ID: 0000-0003-4685-0847}%
}
\author[1,3]{%
	Cong Zhao
}
\author[3]{%
	Wolfgang Krämer
}
\author[1,2]{%
	David Leonhard
}
\author[1,2]{%
	\\Christian Bauckhage
}
\author[1,2]{%
	Rafet Sifa
}
\affil[1]{University of Bonn, Bonn, Germany}
\affil[2]{Fraunhofer IAIS, Sankt Augustin, Germany}
\affil[3]{Deutsche Telekom, Bonn, Germany}

\renewcommand{\shorttitle}{Controlled Randomness Improves the Performance of Transformer Models}

\hypersetup{
pdftitle={Controlled Randomness Improves the Performance of Transformer Models},
pdfsubject={cs.CL, cs.AI},
pdfauthor={Tobias Deu{\ss}er, Cong Zhao, Wolfgang Krämer, David Leonhard, Christian Bauckhage, Rafet Sifa},
pdfkeywords={Natural Language Processing, Regularization, Transformer, Machine Learning},
}

\newcommand\copyrighttext{%
  \footnotesize \textcopyright 2023 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
}
\newcommand\copyrightnotice{%
\begin{tikzpicture}[remember picture,overlay]
\node[anchor=south,yshift=10pt] at (current page.south) {\fbox{\parbox{\dimexpr\textwidth-\fboxsep-\fboxrule\relax}{\copyrighttext}}};
\end{tikzpicture}%
}

\begin{document}
\maketitle
\addtocounter{footnote}{-1}
\begin{abstract}
During the pre-training step of natural language models, the main objective is to learn a general representation of the pre-training dataset, usually requiring large amounts of textual data to capture the complexity and diversity of natural language. Contrasting this, in most cases, the size of the data available to solve the specific downstream task is often dwarfed by the aforementioned pre-training dataset, especially in domains where data is scarce. We introduce controlled randomness, i.e. noise, into the training process to improve fine-tuning language models and explore the performance of targeted noise in addition to the parameters of these models. We find that adding such noise can improve the performance in our two downstream tasks of \textit{joint named entity recognition and relation extraction} and \textit{text summarization}.
\end{abstract}
\copyrightnotice

\keywords{Natural Language Processing \and Regularization \and Transformer \and Machine Learning}

\section{Introduction}

The emergence of pre-trained transformer models brought a massive breakthrough in the field of natural language processing. During pre-training, such transformer models can learn generic language representations with strong generalization capabilities by applying a self-supervised learning approach and leveraging large text corpora. These pre-trained language models can be fine-tuned in various downstream tasks without needing to train from scratch compared to traditional training methods, significantly reducing training costs while achieving excellent performance. Models like BERT \cite{devlin2018bert}, ELECTRA \cite{Clark2020ELECTRA}, or T5 \cite{t5} have achieved remarkable results on several language processing tasks and the most recent developments of even larger language models, made prominent by GPT-3  \cite{gpt3} and GPT-4 \cite{openai2023gpt4} but not limited to these two\footnote{See for example OPT \cite{zhang2022opt}, Bloom \cite{workshop2023bloom}, LLaMA \cite{touvron2023llama}, or Falcon \cite{falcon}.}, improved on these even further. These models have enabled researchers and developers to exploit existing computational linguistic knowledge more conveniently, which in turn has dramatically accelerated the development of natural language processing research and applications.

One of the keys to the success of these models is the ability to adapt to data not encountered during the pre-training phase, i.e. when a downstream task is tackled and the model fine-tuned, as shown in virtually all such applications of transformer models (e.g. \cite{chalkidis2019neural}, \cite{beltagy2020longformer}, \cite{wang2021automated}, \cite{aghajanyan2021muppet}, \cite{kpibert}, \cite{ye2022packed}, \cite{deusser2023contradiction} \cite{ramamurthy2023is}, \cite{deusser2023informed}). To avoid overfitting and instability during this process, one can apply various regularization techniques and data augmentation, which both might help stabilize fine-tuning and improve performance. 

In this work, we build upon the concept introduced in \cite{wu2022noisytune}, namely NoisyTune, which describes the process of adding noise to all the parameters of language models in order to regularize it.

We introduce such noise to more parts of the model to examine how this influences the performance of two downstream tasks, specifically \textit{joint named entity recognition and relation extraction} and \textit{text summarization}. We observe that we can enhance the F$_1$ score of the \textit{joint named entity recognition and relation extraction} model on the KPI-EDGAR dataset \cite{deusser2022kpiedgar} by \textbf{2.977} and the ROUGE--Average\footnote{See subsection~\ref{subsection:data_and_downstream_tasks} for how we define this metric.} score of the \textit{text summarization} model by \textbf{1.387} on the BillSum dataset \cite{kornilova-eidelman-2019-billsum}.

Our contribution is thus investigating how noise added to various parts can improve the performance of the considered downstream task. Therein, we find
that the model's performance is significantly improved by adding controlled noise to certain components. We further theorize that this novel approach can help improve the generalisation of large language models to small datasets or low-resource languages.

In the following, we first review related work with a focus on regularization techniques for natural language processing as well as previous studies on \textit{joint named entity recognition and relation extraction} and \textit{text summarization}. Section~\ref{section:methodology} describes our methodology, i.e., how and where we introduce noise to the model and the general model architecture of both downstream tasks. Thereafter, in Section~\ref{section:experiments}, we outline our dataset, present our experiments, and discuss the results. Section~\ref{section:conclusion} then adds concluding remarks and an outlook into conceivable future work.

\section{Related Work}
In this section, we discuss various other studies conducted on the effect of regularization techniques on machine learning models. Afterwards, we shortly introduce the most relevant advances in both of our tasks, \textit{joint named entity recognition and relation extraction} and \textit{text summarization}.

\subsection{Regularization}

Overfitting and thus regularization of machine learning models have been thoroughly studied since the emergence of the field. Early examples of regularization include the work by \cite{hoerl1970ridge}, \cite{hanson1988comparing}, \cite{mccloskey1989catastrophic}, \cite{breiman1995better}, \cite{girosi1995regularization}, and \cite{prechelt2002early}.
Using noise as a regularization technique for training machine learning models was first considered in \cite{bishop1995training}.

\cite{hinton2012improving} introduced the popular dropout method, which randomly omits a certain part of the feature detectors on each training case. Building on the idea of weight decay \cite{hanson1988comparing}, \cite{loshchilov2018decoupled} investigated the effects of decoupled weight decay regularization on the training of deep neural networks.

Applying such techniques to natural language processing has gained more importance with the ever-increasing 
size of the number of parameters of language models. \cite{merity2018regularizing} considered the
challenge of word-level language modelling and investigated strategies for regularizing Long Short Term Memory-based \cite{lstm} models. \cite{Lee2020Mixout} 
proposed to add dropout to randomly mix pre-trained parameters into the downstream model to reduce forgetting in BERT fine-tuning. Furthermore, \cite{dodge2020finetuning} 
proposed an early stopping method to filter out poor-performing random seeds. On the topic of machine translation, \cite{ott2018analyzing} analyzed uncertainty in machine translation and proposed tools to assess model calibration. \cite{jiang-etal-2020-smart} 
used smoothness-inducing regularization, which tries to effectively manage the complexity of the model to encourage models to be smooth within neighbourhoods of all the inputs.
It was theorized by by \cite{sun2019fine} that smaller learning rates during fine-tuning help the model retain prior knowledge while still adapting to the new task. In \cite{howard-ruder-2018-universal}, the authors 
proposed a gradual unfreezing strategy in which layers of the pre-trained language model are unfrozen one at a time during fine-tuning to cope with catastrophic forgetting. \cite{pan2023improving} 
introduced an extra class-aware initialization stage before fine-tuning and concluded that, in this way, self-supervised models should be easier to train to discriminate between different classes.
Finally, \cite{wu2022noisytune} described how to regularize a language model by adding noise to its weight parameters, which we will use as a further baseline for our approach. By investigating the effect of noise injection into the individual parts of the model explicitly and measuring the performance on two discrete downstream tasks, our work succeeds \cite{wu2022noisytune} significantly by adding specificity to the analysis.

\subsection{Joint Named Entity Recognition and Relation Extraction}

When it comes to the first of the evaluated downstream natural language processing tasks, there is introductory work by \cite{miwa2014modelling},  \cite{li2014incremental}, and \cite{gupta2016table} that was further confirmed by \cite{kambar2022survey}, demonstrating the advantage of joining the subtasks of named entity recognition (NER) and relation extraction (RE) together. Studies, evaluating further concepts to improve the performance of joint NER and RE (also called JNERE), were carried out by \cite{zheng2017joint}, who introduced a novel tagging scheme, \cite{bekoulis2018joint} and \cite{geng2021joint}, who treated the task as a multi-head selection problem, \cite{fu2019graphrel}, who used graph convolutional networks (GCNs), \cite{giorgi2019end} and \cite{xue2019fine}, who leveraged the pre-trained model BERT, \cite{yu2019joint}, who improved performances by introducing a new task decomposition strategy, and \cite{shang2022onerel}, who focused on the problems of cascading errors and redundant information in previous models by creating their own, OneRel, which treats joint extraction as a fine-grained triple classification
problem.

There is also insight on the application of JNERE regarding work by \cite{bhatia2019comprehend}, \cite{chen2020joint}, and \cite{jabbari2020french} for the medical, legal, and financial language domain respectively.

The work by \cite{deusser2022kpiedgar}, which was already mentioned in the introduction, also lays groundwork for the examinations undertaken in this study. However, to the best of our knowledge, the effect of deliberately incorporating noise into the training process of JNERE models specifically has not yet been
systematically
researched.

\subsection{Text Summarization}

As for the task of text summarization, there are several studies conducted on the same dataset as our experiments, that is BillSum, introduced by \cite{kornilova-eidelman-2019-billsum}. Most of these make use of or compare their proposed models to models based on BERT, such as \cite{an2021retrievalsum}, who used a model that incorporates additional knowledge into the summarization task, \cite{abdel2022performance}, who compared various BERT variants, \cite{liang2022efficient}, who proposed the Coarse-to-
Fine Facet-Aware Ranking (C2F-FAR) framework for unsupervised long document summarization, and \cite{jain2022improving}, who leveraged the Kullback-Leibler based summarization. Different but general, model-comparative approaches were, for example, undertaken by \cite{rehman2022analysis} and \cite{mahmoud2022novel}.

BERT was, although on other datasets, also used for text summarization by \cite{zhang2019pretraining}, who used the model for the generation of the abstractive summarization output, and \cite{koniaris2023evaluation}, who worked on greek legal texts. Other legal domain automatic text summarization techniques were studied by, among, but not limited to, \cite{liu2019text}, who proposed a general framework for extractive and abstractive models, and \cite{sheik2021deep}, who focused on improvements through preparation mechanisms and throughout various baselines.

The effects of deliberately introducing noise to the specific task of text summarization were studied by \cite{yousefi2017text}, who presented extractive summarization methods based on term-frequency. \cite{liu2020noisy} used noise to better model uncertainty during training with a student and a teacher model interacting with each other. Both reported improvements using noise at the respective steps in the model training processes. Our work then takes a more fundamental and systematic approach, thoroughly studying the different effects of noise on the different steps of the training process.

\section{Methodology}
\label{section:methodology}

In this section, we first describe how we add noise to the language models to regularize them. Afterwards, we briefly touch upon the models we use for our two tasks, \textit{joint named entity recognition and relation extraction} and \textit{text summarization}.

\subsection{Regularization by adding noise}

The parameter matrices of a language model are denoted as $[\bm{W}_1,\bm{W}_2, ..., \bm{W}_N]$, where $N$ is the number of parameter matrix types. As shown in \cite{wu2022noisytune}, the perturbed version $\tilde{\bm{W}}_i$ of the parameter matrix $\bm{W}_i$ is defined as

\begin{equation}
    \tilde{\bm{W}}_i = \bm{W}_i + \text{U}\left(- \frac{\lambda}{2}, \frac{\lambda}{2} \right) \cdot \sigma \left( \bm{W}_i \right),
    \label{eq:noisytune}
\end{equation}


\noindent where $\text{U}\left(a, b \right)$ represents uniformely distributed noise ranging from $a$ to $b$, $\lambda$ is the fine-tunable hyperparameter that controls the noise intensity, and $\sigma \left( c \right)$ denotes the standard deviation of $c$. Clearly, parameters with a larger variance are subject to stronger noise due to the multiplication of the noise term with the standard deviation $\sigma \left( c \right)$.

In addition to simply adding such noise to all parameters, as seen in \cite{wu2022noisytune}, we investigate how the performance of the downstream task is affected when we only partially inject the noise term into certain parts of the model.
More precisely, we add noise to either the bias term, the weight term, or both with different intensities. Furthermore, during the relation extraction downstream task, we add noise to the residual connection and the layer normalization step. We also divide the BERT \cite{devlin2018bert} encoder into two separate \textit{layer zones} to add noise in different intensities, as \cite{tenney-etal-2019-bert} theorized that BERT solves various language understanding tasks at different layer depths. On the other hand, during the text summarization task, we insert noise separately into the encoder and decoder parts of the model.

Therefore, we adjust Equation~\ref{eq:noisytune} by only perturbing a certain part of the parameter matrix $\bm{W}_i$:

\begin{equation}
    \tilde{\bm{W}}^{\text{loc}}_i = \bm{W}^{\text{loc}}_i + \text{U}\left(- \frac{\lambda}{2}, \frac{\lambda}{2} \right) \cdot \sigma \left( \bm{W}^{\text{loc}}_i \right),
    \label{eq:partual_noise}
\end{equation}

\noindent where $\bm{W}^{\text{loc}}_i$ is the localized, i.e. restricted to certain parts of the model, parameter matrix. Then, $\tilde{\bm{W}}^{\text{loc}}_i$ is the perturbed localized parameter matrix. 

\subsection{Joint named entity recognition and relation extraction}

The \textit{joint named entity recognition and relation extraction} task is defined as extracting entities from a text segment, mostly sentences, and linking them together afterwards. Given this sentence from \cite{deusser2022kpiedgar},

\begin{tcolorbox}[notitle,boxrule=0pt,
boxsep=0pt,left=0.6em,right=0.6em,top=0.5em,bottom=0.5em,
colback=gray!10,
colframe=gray!10]
``In 2021 and 2020 the $\underset{\text{{\color{unibonnblue}{kpi}}}}{\text{{\color{unibonnblue}{total net revenue}}}}$ was \$$\underset{\text{{\color{unibonnblue}{cy}}}}{\text{{\color{unibonnblue}{100}}}}$ million and \$$\underset{\text{{\color{unibonnblue}{py}}}}{\text{{\color{unibonnblue}{80}}}}$ million, respectively.''
\end{tcolorbox}

\noindent one should extract and find the relations

\begin{tcolorbox}[notitle,boxrule=0pt,
boxsep=0pt,left=0.6em,right=0.6em,top=0.5em,bottom=0.5em,
colback=gray!10,
colframe=gray!10]
$\underset{\text{{\color{unibonnblue}{kpi}}}}{\text{{\color{unibonnblue}{total net revenue}}}} - \underset{\text{{\color{unibonnblue}{cy}}}}{\text{{\color{unibonnblue}{100}}}}$,
$\underset{\text{{\color{unibonnblue}{kpi}}}}{\text{{\color{unibonnblue}{total net revenue}}}} - \underset{\text{{\color{unibonnblue}{py}}}}{\text{{\color{unibonnblue}{80}}}}$.
\end{tcolorbox}

To solve the joint named entity recognition and relation extraction task, we employ the model introduced in \cite{kpibert} titled KPI-BERT. Said model has three main building blocks: A BERT-based sentence encoder, a named entity recognition decoder, and a relation extraction decoder. 

To give more detail, given an input sentence tokenized using WordPiece \cite{schuster2012japanese}, we utilize the pre-trained BERT model to obtain the encoded token embeddings. Subsequently, we employ a pooling function to generate word representations by combining the embeddings of individual subwords with a trainable recurrent neural network (RNN) pooling mechanism introduced by \cite{kpibert}. The RNN pooling mechanism is built on a bidirectional gated recurrent unit (GRU) \cite{cho2014properties}. Additionally, we employ conditional label masking to sequentially tag entities before classifying their relations. 

\subsection{Text summarization}

The \textit{text summarization} task involves generating a shorter version of a given text while preserving its vital information. It can be approached through extractive or abstractive methods. Extractive methods involve selecting and combining critical sentences from the original text. 

In contrast, abstractive methods involve generating new sentences that capture the essence of the original text, which is the approach we choose to solve. We pick the Longformer-Encoder-Decoder \cite{beltagy2020longformer} setup and pre-trained model to solve the text summarisation task. It is a Longformer variant with encoder and decoder transformer stacks, but the encoder uses Longformer's efficient local and global attention model instead of the initial fully self-attentive one. Additionally, the decoder uses full self-attention for the encoded token and previously decoded locations.

\section{Experiments}
\label{section:experiments}

Here, we first shed some light on the two datasets for our two tasks, namely \textit{joint named entity recognition and relation extraction} and \textit{text summarization}, and how we evaluated these two. Then, we take a closer look at the results and sum them up. All experiments were conducted on a single Nvidia GeForce RTX 2060 SUPER, and the code is implemented in PyTorch.


We execute the experiments within four phases, starting without noise for the downstream tasks. Second, we add noise to all weights as seen in \cite{wu2022noisytune}. These two approaches can be seen as our two baselines. Finally, we arrived at our results by adding noise only to the bias or weight parts. For joint named entity recognition and relation extraction, noise is additionally added to the residual connections and layer normalization and the layer zones. For text summarization, noise is added separately to the encoder and decoder.

Furthermore, to not cherry-pick any particularly well-performing results, as warned about in \cite{trosten2023questionable}, we run each configuration five times with a different seed each and compare the average of these runs.

\subsection{Data and downstream tasks}
\label{subsection:data_and_downstream_tasks}

The joint named entity recognition and relation extraction, defined in the KPI-EDGAR dataset \cite{deusser2022kpiedgar}, aims to extract information from financial documents, including key performance indicators. It holds a total of 1,355 sentences holding 4,522 entities and 3,841 relations, and is split into a training, validation, and test set, encompassing 969, 146, and 240 sentences each. The named entity recognition part is realized in finding numerical and non-numerical entities, whereas the relation extraction part links the found entities to allow for a meaningful extraction of them. \cite{deusser2022kpiedgar} also defined an \textit{adjusted} F$_1$ score capturing when an entity is only partially found, which we will also be using.

For this \textit{adjusted} F$_1$ score, the \textit{true positives} ($\text{tp}$), \textit{false negatives} ($\text{fn}$), and \textit{false positives} ($\text{fp}$) of a relation $r$ between entity $i$ and $j$ are given by:

\begin{align}
    \text{tp}_r &= \frac{1}{2} \left( \frac{o_i}{n_{i, \text{gt}}} + \frac{o_j}{n_{j, \text{gt}}} \right) \\
    \text{fn}_r &= 1 - \text{tp}_r \\
    \text{fp}_r &= \frac{1}{2} \left( \frac{n_{i, \text{pred}} - o_i}{n_{i, \text{pred}}} +  \frac{n_{j, \text{pred}} - o_j}{n_{j, \text{pred}}} \right), \label{eqf1}
\end{align}

\noindent where 

\begin{align}
    o_i := \left| e_{i, \text{pred}} \cap e_{i, \text{gt}} \right| , \label{eqoverlap}
\end{align}

\noindent is the overlap/intersection $o_i$ of an entity prediction $i$ and its ground truth and $e_{i, \text{pred}}$ and $e_{i, \text{gt}}$ is the set of all token identifiers for the entity prediction and ground truth, respectively. The operation $|\cdot|$ calculates the size of a given set.

For the text summarization task, we consider the BillSum dataset \cite{kornilova-eidelman-2019-billsum}, which aggregates U.S. congressional and California state bills and was split into 18,949 
training bills and 3,269 test bills. We evaluate our results using the ROUGE metrics proposed in \cite{lin-2004-rouge}. Furthermore, to make a convenient comparison possible, we average the values of the ROUGE--1 F$_1$, ROUGE--2 F$_1$, ROUGE--L F$_1$ and ROUGE--L--sum and title it \textit{ROUGE--Average}.

\subsection{Results}

\begin{table}[t]
\centering
    \rowcolors{2}{gray!10}{white}
\begin{tabular}{lcc}
\toprule
\rowcolor{white} Noise added to & $\lambda$ & F$_1$ in  \%
\\
\midrule
Nothing &0 & 40.696
\\

All & 0.81 & 42.824
\\

\textbf{Bias} & \textbf{0.41} & \textbf{43.672}
\\

Weights & 0.50 & 43.084
\\

Add$\&$Norm & 0.2 & 43.411
\\

Layer zones & 0.9 & 42.200
\\
\bottomrule

\end{tabular}


    \caption{Results of adding noise to KPI-BERT \cite{kpibert}, evaluated on KPI-EDGAR \cite{deusser2022kpiedgar}.  Adding noise to all parameters is equivalent to the approach from \cite{wu2022noisytune}. Add\&Norm 
    refers to the process of adding noise to residual connections and layer normalization.}
    \label{tab:kpiedgar_main}
\end{table}

\begin{table}[t]
\centering
    \rowcolors{2}{gray!10}{white}
\begin{tabular}{lcc}
\toprule
\rowcolor{white} Noise added to & $\lambda$ & ROUGE--Average in  \%
\\
\midrule
Nothing &0 &  34.148
\\

All & 0.3 & 35.131
\\

Bias & 0.4 &  34.854
\\

Weights & 0.1 &  34.703
\\

Encoder & 0.1 &  34.709
\\

\textbf{Decoder} & \textbf{0.8} & \textbf{35.534}
\\
\bottomrule

\end{tabular}
    \caption{Results of adding noise to the Longformer-Encoder-Decoder \cite{beltagy2020longformer}, evaluated on BillSum \cite{kornilova-eidelman-2019-billsum}.  Adding noise to all parameters is equivalent to the approach from \cite{wu2022noisytune}.}
    \label{tab:billsum_main}
\end{table}

As seen in Table~\ref{tab:kpiedgar_main} and \ref{tab:billsum_main}, adding noise, in general, does help the models generalize better. Thus, we can confirm the findings in \cite{wu2022noisytune}. Additionally, we are able to improve upon that by exposing only certain parts of the models to noise.

Interestingly, there is no simple \say{best approach} on where to add noise in a model, as seen in the different best-performing noise locations of Table~\ref{tab:kpiedgar_main} and \ref{tab:billsum_main}. Therefore, this is another hyperparameter that has to be fine-tuned. In \cite{wu2022noisytune}, they only tested a limited range, i.e. $\lambda \in \{0, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3\}$, and they found that $0.1 \leq \lambda \leq 0.15$ were optimal. Contrasting this, in our experiments, the best results are obtained with noise intensities significantly larger than this, as demonstrated in Table~\ref{tab:kpiedgar_main} and \ref{tab:billsum_main}.

Still, we found that our approach boosted both performances. In the case of KPI-BERT \cite{kpibert} on KPI-EDGAR \cite{deusser2022kpiedgar}, we achieve a remarkable increase of \textbf{2.977}\% F$_1$ compared to the non-perturbed model and an increase of \textbf{0.849}\% F$_1$ compared to the NoisyTune approach \cite{wu2022noisytune}. On BillSum \cite{kornilova-eidelman-2019-billsum}, our approach improved the performance of the Longformer-Encoder-Decoder \cite{beltagy2020longformer} by \textbf{1.387}\% and \textbf{0.403}\% compared to the non-perturbed model and the NoisyTune baseline, respectively.

\section{Conclusion and Future Work}
\label{section:conclusion}

In this paper, we study the effect of controlled randomness on transformer models and how such noise, introduced into various parts of the model as shown in Equation~\ref{eq:partual_noise}, can be seen as a potent regularization tool for ever-increasing language models. We studied two different downstream tasks, namely \textit{joint named entity recognition and relation extraction} and \textit{text summarization}, and found that if certain parts of their respective transformer models are infused with a noise component, we can increase their performance significantly. In the first task, three out of the four scenarios, i.e. adding noise to either the bias term, the weights, or to residual connections and layer normalization, were successful and demonstrated additional enhancements over our already string baseline introduced in \cite{wu2022noisytune}. In the second task, we can still improve on this baseline, but only one of our scenarios offers an improvement, namely adding noise the whole decoder and leaving the encoder as it is.

There are two further conclusions from our results. One is that the optimal value of the hyperparameter $\lambda$, as defined in Equation~\eqref{eq:partual_noise}, depends heavily on the dataset and on the task. The other is that in the paper introducing our baseline NoisyTune \cite{wu2022noisytune}, they only tested the range $\lambda \in \{0, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3\}$, which should have been extended further to achieve even better results, as seen in Table~\ref{tab:kpiedgar_main} and \ref{tab:billsum_main}.

In future work, we plan to tackle even more downstream tasks, like named entity recognition, natural language inference, or sentiment analysis, to see if we can achieve such positive results on these natural language processing applications as well. One could also apply this regularization technique to any pre-trained large language model like Llama \cite{touvron2023llama} or Bloom \cite{workshop2023bloom} and experiment on how these can then handle themselves. It might partly alleviate the flaw of requiring large datasets for fine-tuning  when training data is sparse. 

Another interesting idea would be trying out different noise distributions than the uniformly distributed one seen in Equation~\eqref{eq:partual_noise}. One candidate for this could be a Gaussian distribution with fat tails, as this does not have such extreme cut-off points as a uniform distribution.

Furthermore, it would be immensely interesting to see if some low-resource languages can benefit from such a regularization approach. We theorize that in this case, one might see even larger increases in performance, as the \say{seen language} imbalance of multilingual models, i.e. the distribution of different languages in the training dataset, is usually quite pronounced and heavily skewed towards English, as seen in e.g. \cite{conneau-etal-2020-unsupervised} or \cite{laurenccon2022bigscience}.

\section*{Acknowledgments}

This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence. 

\bibliographystyle{plainnat}
\bibliography{references}

\end{document}